# FDDM: Unsupervised Medical Image Translation with a Frequency-Decoupled Diffusion Model

### v2

@@ 3.2.2 Pelvis Dataset
Step $\tilde{T}$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
coarse CT 39.8187 0.9123 $\pm$ 0.0386 37.7208 $\pm$ 2.7626
without IPM 26.6595 0.8973 $\pm$ 0.0467 37.7703 $\pm$ 3.1372
without EPM 27.9621 0.9119 $\pm$ 0.0390 38.0098 $\pm$ 2.8670
FDDM 25.8565 0.9144 $\pm$ 0.0379 38.0815 $\pm$ 2.8457
@@ 3.3 Ablation Study
Step $\tilde{T}$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
100 28.2487 0.9166 $\pm$ 0.0372 38.0291 $\pm$ 2.8408
300 25.5146 0.9022 $\pm$ 0.0433 37.7681 $\pm$ 2.9515
600 27.1035 0.8896 $\pm$ 0.0476 37.5074 $\pm$ 2.9805
900 37.3593 0.7926 $\pm$ 0.0904 35.8834 $\pm$ 2.3607
@@
Additionally, our diffusion model employs a dual-path reverse diffusion process, using an implicit probabilistic model without random noise for low frequencies and an explicit probabilistic model with random noise for high frequencies. The relevant results are shown in Table 4. When the implicit probabilistic model is removed, leaving only the explicit probabilistic model with random noise, SSIM and PSNR decrease while FID is more stable, indicating a decline in anatomical accuracy. It suggests that removing the implicit probabilistic model primarily affects anatomical accuracy as using only the explicit probabilistic model introduces additional randomness. However, when the explicit probabilistic model is removed, FID increases more, indicating a deterioration in image quality. During the reverse diffusion process, the model transitions from step $t$ to $t-1$ by predicting noise to obtain a clean CT and then adding the predicted noise with a smaller coefficient. This forms the basis of the diffusion model, allowing each diffusion step to remain stable. However, typically in explicit probabilistic models (such as DDPM), random Gaussian noise is also added at each step as a perturbation term to disrupt the existing image. By perturbing the existing information, the model can generate new information closer to the CT. Thus when the explicit probabilistic model is removed, the perturbation power of the diffusion model is reduced which degrades the FID. The same observation can be made in Fig. 7. In the first example of Fig. 7, the coarse CT has erroneous anatomical structures around the eyes. The results without the explicit probabilistic model are very close to the coarse CT, indicating that the model did not correct the errors. In contrast, the results with the explicit probabilistic model corrected the errors in the coarse CT, making the anatomical structures closer to the ground truth. However, the second example in Fig. 7 shows that using only the explicit probabilistic model can introduce excessive randomness, leading to erroneous generation. Therefore, our dual-path reverse diffusion aims to strike a balance, correcting the errors in the coarse CT while preventing the model from generating overly randomized results.
3.4 Impact of the Forward Diffusion Step $T_{s}$
@@
For FDDM, although the diffusion model is trained with 1000 forward and reverse diffusion steps, we can customize this number of steps during testing which allows the diffusion model to act upon varying degrees of low-frequency information. The forward diffusion process approximates a low-pass filter, and the more steps, the less image information is retained. If the number of forward diffusion steps is too small, errors from the coarse CT image obtained by VAE conversion will be overly retained , affecting the final CT accuracy. However, if the number of steps is too large, too little image information is retained to effectively guide the reverse diffusion process. Therefore, we need to find an optimal $T_{s}$. In Table 5 and Fig. 8, we tested the model’s performance with different $T_{s}$. After comparison, we chose $T_{s}=300$ for our experimental setting. To minimize the compounding factors, we did not include CT boundary guidance information when testing the performance of different $T_{s}$. By comparing the results of $T_{s}=300$ in Table 5 and Table 4, we found that the FID score is slightly better without CT boundary guidance, but at the cost of a significant decrease in SSIM and PSNR when $T_{s}$ is large. The lower FID with CT boundary information-based guidance is because the CT boundary information is generated by the initial conversion module, which contains imperfections slightly affecting the image quality of final output CT. However, the high-frequency CT boundary information serves as strong guidance to preserve the anatomical accuracy as demonstrated in the previous work [29]. As shown in Table 5, without the CT boundary guidance the structure integrity deteriorates quickly when the high-frequency information is removed with relatively large $T_{s}$ values.
4 Conclusion and Future Work

### v3

@@ IV.B.2. Pelvis Dataset
Step $\tilde{T}$ FID $\downarrow$ SSIM $\uparrow$ MSE (HU) $\downarrow$ MAE (HU) $\downarrow$
coarse CT 39.8187 0.8916 $\pm$ 0.0534 14994.927 $\pm$ 8366.118 35.2006 $\pm$ 14.8481
without IPM 26.6595 0.8839 $\pm$ 0.0558 14553.736 $\pm$ 7875.460 32.8674 $\pm$ 14.4395
without EPM 27.9621 0.9020 $\pm$ 0.0462 13881.806 $\pm$ 7730.327 31.5993 $\pm$ 13.2413
FDDM 25.8565 0.9116 $\pm$ 0.0409 11862.303 $\pm$ 6785.043 27.8005 $\pm$ 12.3200
@@ IV.C. Ablation Study
Step $\tilde{T}$ FID $\downarrow$ SSIM $\uparrow$ MSE (HU) $\downarrow$ MAE (HU) $\downarrow$
100 28.2487 0.9136 $\pm$ 0.0401 11748.454 $\pm$ 6820.354 27.6307 $\pm$ 12.3094
300 25.5146 0.9089 $\pm$ 0.0417 12345.224 $\pm$ 7052.862 28.5101 $\pm$ 12.6511
600 27.1035 0.8808 $\pm$ 0.0527 17777.790 $\pm$ 9870.868 36.0422 $\pm$ 16.1415
900 37.3593 0.7667 $\pm$ 0.0965 54486.934 $\pm$ 25457.920 78.4379 $\pm$ 31.5931
@@
Additionally, our diffusion model employs a dual-path reverse diffusion process, using an implicit probabilistic model without random noise for low frequencies and an explicit probabilistic model with random noise for high frequencies. The relevant results are shown in Table 4. When the implicit probabilistic model is removed, leaving only the explicit probabilistic model with random noise, SSIM, MSE, and MAE decrease while FID is more stable, indicating a decline in anatomical accuracy. It suggests that removing the implicit probabilistic model primarily affects anatomical accuracy as using only the explicit probabilistic model introduces additional randomness. However, when the explicit probabilistic model is removed, FID increases more, indicating a deterioration in image quality. During the reverse diffusion process, the model transitions from step $t$ to $t-1$ by predicting noise to obtain a clean CT and then adding the predicted noise with a smaller coefficient. This forms the basis of the diffusion model, allowing each diffusion step to remain stable. However, typically in explicit probabilistic models (such as DDPM), random Gaussian noise is also added at each step as a perturbation term to disrupt the existing image. By perturbing the existing information, the model can generate new information closer to the CT. Thus when the explicit probabilistic model is removed, the perturbation power of the diffusion model is reduced which degrades the FID. The same observation can be made in Fig. 7. In the first example of Fig. 7, the coarse CT has erroneous anatomical structures around the eyes. The results without the explicit probabilistic model are very close to the coarse CT, indicating that the model did not correct the errors. In contrast, the results with the explicit probabilistic model corrected the errors in the coarse CT, making the anatomical structures closer to the ground truth. However, the second example in Fig. 7 shows that using only the explicit probabilistic model can introduce excessive randomness, leading to erroneous generation. Therefore, our dual-path reverse diffusion aims to strike a balance, correcting the errors in the coarse CT while preventing the model from generating overly randomized results.
IV.D. Impact of the Forward Diffusion Step $T_{s}$
@@
For FDDM, although the diffusion model is trained with 1000 forward and reverse diffusion steps, we can customize this number of steps during testing which allows the diffusion model to act upon varying degrees of low-frequency information. The forward diffusion process approximates a low-pass filter, and the more steps, the less image information is retained. If the number of forward diffusion steps is too small, errors from the coarse CT image obtained by VAE conversion will be overly retained , affecting the final CT accuracy. However, if the number of steps is too large, too little image information is retained to effectively guide the reverse diffusion process. Therefore, we need to find an optimal $T_{s}$. In Table 5 and Fig. 8, we tested the model’s performance with different $T_{s}$. After comparison, we chose $T_{s}=300$ for our experimental setting. To minimize the compounding factors, we did not include CT boundary guidance information when testing the performance of different $T_{s}$. By comparing the results of $T_{s}=300$ in Table 5 and Table 4, we found that the FID score is slightly better without CT boundary guidance, but at the cost of a significant decrease in SSIM, MSE, and MAE when $T_{s}$ is large. The lower FID with CT boundary information-based guidance is because the CT boundary information is generated by the initial conversion module, which contains imperfections slightly affecting the image quality of final output CT. However, the high-frequency CT boundary information serves as strong guidance to preserve the anatomical accuracy as demonstrated in the previous work [29]. As shown in Table 5, without the CT boundary guidance the structure integrity deteriorates quickly when the high-frequency information is removed with relatively large $T_{s}$ values.
V. Conclusion and Future Work
@@
  • [1] S. Kazemifar, S. McGuire, R. Timmerman, Z. Wardak, D. Nguyen, Y. Park, S. Jiang, and A. Owrangi, “Mri-only brain radiotherapy: Assessing the dosimetric accuracy of synthetic ct images generated using a deep learning approach,” Radiotherapy and Oncology, vol. 136, pp. 56–63, 2019.
  • [2] G. Payne and M. Leach, “Applications of magnetic resonance spectroscopy in radiotherapy treatment planning,” The British journal of radiology, vol. 79, no. special_issue_1, pp. S16–S26, 2006.
  • [3] A. Thornton Jr, H. Sandler, R. Ten Haken, D. McShan, B. Fraass, M. LaVigne, and B. Yanks, “The clinical utility of magnetic resonance imaging in 3-dimensional treatment planning of brain neoplasms,” International Journal of Radiation Oncology* Biology* Physics, vol. 24, no. 4, pp. 767–775, 1992.
  • [4] M. Goitein, J. Wittenberg, M. Mendiondo, J. Doucette, C. Friedberg, J. Ferrucci, L. Gunderson, R. Linggood, W. U. Shipley, and H. V. Fineberg, “The value of ct scanning in radiation therapy treatment planning: a prospective study,” International Journal of Radiation Oncology* Biology* Physics, vol. 5, no. 10, pp. 1787–1798, 1979.
  • [5] C. B. Saw, A. Loper, K. Komanduri, T. Combine, S. Huq, and C. Scicutella, “Determination of ct-to-density conversion relationship for image-based treatment planning systems,” Medical Dosimetry, vol. 30, no. 3, pp. 145–148, 2005.
  • [6] H.-C. Shao, Y. Li, J. Wang, S. Jiang, and Y. Zhang, “Real-time liver tumor localization via combined surface imaging and a single x-ray projection,” Physics in Medicine & Biology, vol. 68, no. 6, p. 065002, 2023.
  • [7] Y. Zhang, F. Ye, L. Chen, F. Xu, X. Chen, H. Wu, M. Cao, Y. Li, Y. Wang, and X. Huang, “Children’s dental panoramic radiographs dataset for caries segmentation and dental disease detection,” Scientific Data, vol. 10, no. 1, p. 380, 2023.
  • [8] H.-C. Shao, Y. Li, J. Wang, S. Jiang, and Y. Zhang, “Real-time liver motion estimation via deep learning-based angle-agnostic x-ray imaging,” Medical physics.
  • [9] H. Chandarana, H. Wang, R. Tijssen, and I. J. Das, “Emerging role of mri in radiation therapy,” Journal of Magnetic Resonance Imaging, vol. 48, no. 6, pp. 1468–1478, 2018.
  • [10] J. M. Edmund and T. Nyholm, “A review of substitute ct generation for mri-only radiation therapy,” Radiation Oncology, vol. 12, pp. 1–15, 2017.
  • [11] K. Wang, Y. Li, M. Dohopolski, T. Peng, W. Lu, Y. Zhang, and J. Wang, “Recurrence-free survival prediction under the guidance of automatic gross tumor volume segmentation for head and neck cancers,” in 3D Head and Neck Tumor Segmentation in PET/CT Challenge, pp. 144–153, Springer, 2022.
  • [12] I. Goodfellow, J. Pouget-Abadie, M. Mirza, B. Xu, D. Warde-Farley, S. Ozair, A. Courville, and Y. Bengio, “Generative adversarial networks,” Communications of the ACM, vol. 63, no. 11, pp. 139–144, 2020.
  • [13] D. P. Kingma, M. Welling, et al., “An introduction to variational autoencoders,” Foundations and Trends® in Machine Learning, vol. 12, no. 4, pp. 307–392, 2019.
  • [14] M. Heusel, H. Ramsauer, T. Unterthiner, B. Nessler, and S. Hochreiter, “Gans trained by a two time-scale update rule converge to a local nash equilibrium,” Advances in neural information processing systems, vol. 30, 2017.
  • [15] P. Dhariwal and A. Nichol, “Diffusion models beat gans on image synthesis,” Advances in Neural Information Processing Systems, vol. 34, pp. 8780–8794, 2021.
  • [16] L. Yang, Z. Zhang, Y. Song, S. Hong, R. Xu, Y. Zhao, Y. Shao, W. Zhang, B. Cui, and M.-H. Yang, “Diffusion models: A comprehensive survey of methods and applications,” arXiv preprint arXiv:2209.00796, 2022.
  • [17] Y. Song, J. Sohl-Dickstein, D. P. Kingma, A. Kumar, S. Ermon, and B. Poole, “Score-based generative modeling through stochastic differential equations,” in International Conference on Learning Representations, 2021.
  • [18] J. Song, C. Meng, and S. Ermon, “Denoising diffusion implicit models,” in International Conference on Learning Representations, 2021.
  • [19] J. Ho, A. Jain, and P. Abbeel, “Denoising diffusion probabilistic models,” Advances in Neural Information Processing Systems, vol. 33, pp. 6840–6851, 2020.
  • [20] K. Armanious, C. Jiang, M. Fischer, T. Küstner, T. Hepp, K. Nikolaou, S. Gatidis, and B. Yang, “Medgan: Medical image translation using gans,” Computerized medical imaging and graphics, vol. 79, p. 101684, 2020.
  • [21] S. Kaji and S. Kida, “Overview of image-to-image translation by use of deep neural networks: denoising, super-resolution, modality conversion, and reconstruction in medical imaging,” Radiological physics and technology, vol. 12, no. 3, pp. 235–248, 2019.
  • [22] J.-Y. Zhu, T. Park, P. Isola, and A. A. Efros, “Unpaired image-to-image translation using cycle-consistent adversarial networks,” in Proceedings of the IEEE international conference on computer vision, pp. 2223–2232, 2017.
  • [23] H. Fu, M. Gong, C. Wang, K. Batmanghelich, K. Zhang, and D. Tao, “Geometry-consistent generative adversarial networks for one-sided unsupervised domain mapping,” in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 2427–2436, 2019.
  • [24] L. Kong, C. Lian, D. Huang, Y. Hu, Q. Zhou, et al., “Breaking the dilemma of medical image-to-image translation,” Advances in Neural Information Processing Systems, vol. 34, pp. 1964–1978, 2021.
  • [25] M.-Y. Liu, T. Breuel, and J. Kautz, “Unsupervised image-to-image translation networks,” Advances in neural information processing systems, vol. 30, 2017.
  • [26] X. Huang, M.-Y. Liu, S. Belongie, and J. Kautz, “Multimodal unsupervised image-to-image translation,” in Proceedings of the European conference on computer vision (ECCV), pp. 172–189, 2018.
  • [27] C. Meng, Y. Song, J. Song, J. Wu, J.-Y. Zhu, and S. Ermon, “Sdedit: Image synthesis and editing with stochastic differential equations,” arXiv preprint arXiv:2108.01073, 2021.
  • [28] M. Özbey, O. Dalmaz, S. U. Dar, H. A. Bedel, Ş. Özturk, A. Güngör, and T. Çukur, “Unsupervised medical image translation with adversarial diffusion models,” IEEE Transactions on Medical Imaging, 2023.
  • [29] Y. Li, H.-C. Shao, X. Liang, L. Chen, R. Li, S. Jiang, J. Wang, and Y. Zhang, “Zero-shot medical image translation via frequency-guided diffusion models,” IEEE transactions on medical imaging, 2023.
  • [30] J. Sohl-Dickstein, E. Weiss, N. Maheswaranathan, and S. Ganguli, “Deep unsupervised learning using nonequilibrium thermodynamics,” in International Conference on Machine Learning, pp. 2256–2265, PMLR, 2015.
  • [31] T. Karras, M. Aittala, J. Lehtinen, J. Hellsten, T. Aila, and S. Laine, “Analyzing and improving the training dynamics of diffusion models,” in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 24174–24184, 2024.
  • [32] Z. Wang, Z. Cui, and Y. Zhu, “Multi-modal medical image fusion by laplacian pyramid and adaptive sparse representation,” Computers in Biology and Medicine, vol. 123, p. 103823, 2020.
  • [33] A. Thummerer, E. van der Bijl, A. Galapon Jr, J. J. Verhoeff, J. A. Langendijk, S. Both, C. N. A. van den Berg, and M. Maspero, “Synthrad2023 grand challenge dataset: Generating synthetic ct for radiotherapy,” Medical physics, vol. 50, no. 7, pp. 4664–4674, 2023.
  • [34] E. Huijben, M. L. Terpstra, A. Galapon Jr, S. Pai, A. Thummerer, P. Koopmans, M. Afonso, M. van Eijnatten, O. Gurney-Champion, Z. Chen, et al., “Generating synthetic computed tomography for radiotherapy: Synthrad2023 challenge report,” arXiv preprint arXiv:2403.08447, 2024.